1. Find a basis for the subspace of R4 spanned by the given vectors.

(1,1,-5,-6), (2,0,2,-2), (3,-1,0,8).

2. Find a subset of the vectors that forms a basis for the space spanned by the vectors, then express each vector that is not in the basis as a linear combination of the basis vectors.

v1=(1,0,1,1), v2 = (-4,4,-1,4), v3 = (2,4,5,10), v4 = (-10,4,-7,-2)

3. Find a subset of the vectors that forms a basis for the space spanned by the vectors, then express each vector that is not in the basis as a linear combination of the basis vectors.

v1=(1,-1,5,2), v2 = (-2,3,1,0), v3 = (5,-6,14,6), v4 = (0,4,2,-3), v5 = (3,15,45,2)

4. Determine whether b is in the column space of A, and if so, express b as a linear combination of the column vectors of A.

A = 3 4 B = -1
6 -9 15
1
Expert's answer
2020-01-07T07:45:17-0500

1.As the set { "v_1,v_2,v_3" } span a subspace of "R^{4}"

then, "c_1v_1+c_2v_2+c_3v_3=0"

"c_1(1,1,-5,-6)+c_2(2,0,2,-2)+c_3(3,-1,0,8)=0"

We get following equations from above equation

"c_1+2c_2+3c_3=0"

"c_1-c_3=0"

"-5c_1+2c_2=0"

"-6c_1-2c_2+8c_3=0"

From the above equation we can easily conclude that "c_1=c_2=c_3=0"

So,{ "v_1,v_2,v_3" } are linearly independent.

Thus, the set is the basis for the subspace of R4.

2.As the set "v_1,v_2,v_3,v_4" span R4.

then,"c_1v_1+c_2v_2+c_3v_3+c_4v_4=0"

"c_1(1,0,1,1)+c_2(-4,4,-1,4)+c_3(2,4,5,10)+c_4(-10,4,-7,-2)=0"

We get equations:

"c_1\t-4 c_2\t +2 c_3\t-10 c_4\t=\t0"

"4 c_2\t +4 c_3\t +4 c_4\t=\t0"

"c_1\t-c_2\t +5 c_3\t-7 c_4\t=\t0"

"c_1\t +4 c_2\t +10 c_3\t-2 c_4\t=\t0"

Solving these equations we can easily get

"c_1\t=\t-6 c_3\t+6 c_4"

"c_2\t=\t-c_3\t-c_4"

"c_3\t=\tarbitrary"

"c_4\t=\tarbitrary"

{ "v_1,v_2" } span { "v_3,v_4" }

{"v_1,v_2" } spans a subspace of R4 and vectors are linearly independent.Thus,vectors serve as the basis of a subspace of R4.

3."c_1v_1+c_2v_2+c_3v_3+c_4v_4+c_5v_5=0"

"c_1(1,-1,5,2)+c_2(-2,3,1,0)+c_3(5,-6,14,6)+c_4(0,4,2,-3)+c_5(3,15,45,2)=0"

We get the following equations from the above equation.

"c_1\t-2 c_2\t +5 c_3\t+3 c_5\t=\t0"

"- c_1\t +3 c_2\t-6 c_3\t +4 c_4\t +15 c_5\t=\t0"

"5 c_1\t +c_2\t +14 c_3\t +2 c_4\t +45 c_5\t=\t0"

"2 c_1+6 c_3\t-3 c_4\t +2 c_5\t=\t0"

Solving these equations we can get:

"c_1\t=\t-3 c_3\t-7 c_5"

"c_2\t=c_3\t-2 c_5"

"c_3\t=\tarbitrary"

"c_4\t=\t-4 c_5"

"c_5\t=\tarbitrary"

{"v_1,v_2,v_4" } span {"v_3,v_5" }

{ "v_1,v_2,v_4" } span a subspace of R4 and vectors are linearly independent.Thus, vectors serve as the basis of a subspace of R4.

4."A=\begin{bmatrix}
 3 & 4 \\
 6 & -9
\end{bmatrix}"

"B=\begin{bmatrix}
 -1 \\
 15
\end{bmatrix}"

"B=c_1\begin{pmatrix}
 3 \\
 6
\end{pmatrix}+c_2\begin{pmatrix}
 4 \\
 -9
\end{pmatrix}"


"3c_1+4c_2=-1"

"6c_1-9c_2=15"

"c_1=1,c_2=-1"

Yes, B is in the column space of A as it can be expressed as a linear combination of column vectors of A.